OpenStudy (anonymous):

Simplify (m+11)+(n+44)

pooja195 (pooja195):

Combine Like Terms: m+11+n+44 (m)+(n)+(11+44)=?

OpenStudy (brian22):

m+n+55

OpenStudy (anonymous):

still don't get it

pooja195 (pooja195):

Its just combining like terms

OpenStudy (anonymous):

i got 55 as the answer

OpenStudy (brian22):

go to mathway.com and it give you step by step

pooja195 (pooja195):

1st remove the parentheses :) \[\huge~\rm~m+11+n+44\] 2nd combine like terms \[\huge~\rm~(m)+(n)+(11+44)=m+n+55\]

pooja195 (pooja195):

Thats all it can be simplified to

pooja195 (pooja195):

Or you can write it as mn+55